A chemist wants to make 54 mL of a 17% acid solution by mixing a 13% acid solution and an 19% acid solution. How many milliliters of each solution should the chemist use?
13% solution ? mL
19% solution ? mL

Answer by checkley77(12844)

,19X+.13(54-X)=.17*54
.19X+7.02-.13X=9.18
.06X=9.18-7.02
.06X=2.16
X=2.16/.06
X=36 mL OF THE 19% SOLUTION IS USED.
54-36=18 mL OF THE 13 % SOLUTION IS USED.
PROOF:
.19*36+.13*18=.17*54
6.84+2.34=9.18
9.18=9.18